Kate is 20 years older than meg. In 9 years kate will be twice as old as meg. How old is kate?
Answer by idefk69(1)

M~6+9=15, K~26+9=35
M~7+9=16, K~27+9=36
M~8+9=17, K~28+9=37
M~9+9=18, K~29+9=38
M~10+9=19, K~30+9=39
M~11+9=20, K~31+9=40 (i used the guess and check method) 31 is 20 more than 11,and the 9 is well the 9 years and 40 is double 20. So Kate is 31.